OpenStudy (anonymous):

In order to prototype a two dimensional game, you first need to create a cardboard mock-up. Since you’ll be creating several versions of this game, you want to save money. Therefore, you have to minimize the amount of cardboard used. The play area of the prototype must be 36 square inches. As shown below, the bottom 1.5 inches of the mock-up is reserved for buttons and controls and the right-hand 2 inches is reserved for information and statistics. What are the total dimensions of the mock-up such that the amount of cardboard used is minimized?

OpenStudy (goformit100):

"Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Here is what I have so far I believe I may be off on the domain a. Primary equation Total Area w + 2 * h + 1.5 b. Secondary equation The play area w * h = 36 c. Function of one variable w * h / h = 36 / h w = 36 / h 36 / h + 2 * h + 1.5 d. Feasible domain (0, 36) or (0, 18)

OpenStudy (compassionate):

@goformit100 , @robtobey , @Hero , @AravindG Can any of you help him? I haven't taken Domains.

OpenStudy (anonymous):

I believe this would relate to extrema min/max